OpenStudy (anonymous):

To paint his house, Samuel bought 2 cans of paint and one roller for 30$. When friends came over to help, he went back to the store and bought 3 more cans of paint and 2 more rollers for 50$. How much did he pay for each roller?

OpenStudy (anonymous):

paint -- p roller -- r let's set up a system of equation ------------------------------------------------ 2 can of paint and one roller cost 30 dollar 2p+r=30 3 can of paint and 2 roller cost 50 dollar 3p+2r=50 lets solve that equation 2p+r=30 3p+2r=50 -4p-2r=-60 3p+2r=50 --------------- -p =-10 p =10 2p+r=30 20+r=30 r=10 paint can cost 10 dollar and roller cost 10 dollar as well